# Inference of time delay in stochastic systems

### v1

@@ Abstract
Time delay is ubiquitous in many experimental and real-world situations. It is often unclear whether time delay plays a significant role in observed phenomena, and if it does, how long the time lag really is. This would be invaluable knowledge when analyzing and modeling such systems. Hitherto, no universal method is available by which the time delay can be inferred. To address this problem, we propose and demonstrate two different methods to infer time delay in overdamped Langevin systems with delayed feedback. In the first part, we focus on the power spectral density based on the positional data and use a characteristic signature of the time delay to infer the delay time. In limiting cases, we establish a direct relation of the observations made for nonlinear time-delayed feedback forces to analytical results obtained for the linear system. In other situations despite the absence of this direct relation, the characteristic signature remains and can be exploited by a semiautomatic method to infer the delay time. Furthermore, it may not always desirable or possible to observe a system for a long time to infer dependencies and parameters. Thus, in the second part, we propose for the first time a probing method combined with a neural network to infer the delay time, which requires only short observation time series. These proposed methods for inferring time delays in stochastic systems may prove to be valuable tools for gaining deeper insight into the role of delay across a wide range of applications — from the behavior of individual colloidal particles under feedback control to emergent collective phenomena such as flocking and swarming.
@@ I Introduction
Time delay is ubiquitous in many experimental and real-world situations [1]. In some cases, it is introduced deliberately, whereas in other situations its presence is unavoidable. A prominent example for deliberate use of delay is Pyragas time-delayed feedback for the stabilization of otherwise unstable orbits [2, 3], later applied to stabilization of satellite motion [4]. In experiments involving feedback control, researchers have explicitly considered the time delay during modeling experiments [5, 6, 7]. In colloidal systems exhibiting stochastic dynamics, time delayed feedback control is employed to modify the transport properties [8, 9]. Moveover, delay has been shown to have tremendous effect on thermodynamic observables  [10, 11, 12, 13, 14], as demonstrated in recent experiments [7, 15].
@@
Another area where delay is important and, at the same time difficult to control, are systems with finite-time information propagation. Typical examples include communication lag while steering far objects such as satellites, the reaction time humans experience when reacting to an external stimulus or the time lag in experiments between processing image data and applying the desired steering or control impulse [16]. Moreover, time delay may have to be taken into consideration in complex systems such as, in swarms [17], in financial markets [18, 19], in robotic systems [20], and in traffic flow [21, 22].
@@
Indeed, it has recently been shown in theoretical studies as well as simulations that the introduction of time delay and/or time-delayed feedback can induce intriguing emergent collective behavior. Examples of such systems range from polar active matter with included delayed interactions [23, 24], and Gaussian time-delayed feedback applied to Brownian particles [25, 26, 27], to ecological systems [28], and oscillator networks [29, 30]. However, in many experimental and real-world situations, it is not a priori clear whether time delay plays a significant role [31, 16] or if it is present at all. Therefore, it is of fundamental importance to investigate the presence of time delay based on observable quantities to accurately design and understand physical systems.
@@
In recent years, ideas of parameter inference for dynamical systems and equation discovery (for ordinary as well as partial differential equations) from (noisy) data have received considerable interest. Methods range from Bayesian inference [32] to the application of neural network-based techniques [33, 34]. Although for some types of system, methods for inferring parameters for delay differential equations (DDE) [35] and specifically time delay from time series data [36, 37, 38] exist, those methods are typically designed for application to deterministic systems or noisy measurement data [37]. For fully stochastic systems, on the other hand, to the best of our knowledge, no universal approach for the inference of time delay exists. Motivated by this gap, we here for the first time propose and demonstrate two approaches to infer time delay for Brownian systems. We thereby focus on systems with a specific delay time rather than distributed delay.
@@
The first approach is based on the power spectral density (PSD) of stochastic time series, which is often used to analyze experimental data [39, 40, 41] including systems with anomalous diffusion [42]. Here, we identify a characteristic signature of time delay, namely oscillations with a delay-specific period and demonstrate in a proof of concept how this signature can be exploited to infer the delay time. To do this we start from Brownian systems subject to linear time-delayed feedback and then consider highly nonlinear time-delayed feedback. We thereby demonstrate that the characteristic signature carries over from systems subject to linear feedback forces to nonlinear systems. This makes the PSD-based inference approach agnostic with respect to the functional form of the underlying stochastic equations of motion. The method is not limited to stationary states in stochastic systems, but can be extended beyond non-equilibrium steady states to non-stationary problems.
@@
The PSD approach, however, has a drawback. A large number (typically $N\gtrsim 100$) of relatively long [several Brownian times $\tau_{B}$ (defined in Sec. II)] trajectories are required to infer the time delay with sufficient precision. In real-world applications, having to record a large number of long trajectories may (depending on the scenario) not be desirable or not even possible.
@@
Given the potential difficulties, we propose a powerful alternative approach employing a neural network. To establish the concept, we utilize the fact that systems that involve time delay, rather than instantaneously responding to an external stimulus, react with some delay. Therefore, by applying a suitable perturbation and observing the dynamical response, we can obtain information about the underlying delay time of the system. In order to classify the system’s response, based on trajectory time series data, we make use of a convolutional neural network [43, 44]. Such a network can be trained to output the most likely time delay for one or even several different functional forms of the underlying delayed and instantaneous forces. Compared to the PSD approach, the main advantage of this approach is that the observation window (i.e., the trajectory length) only has to be of the order of the time delay itself, and relatively few trajectories (typically $N\sim 25$) are sufficient for precise inference. In combination with state-of-the-art computer systems, this brings the inference of time delay in real time within the realm of possibilities. Furthermore, since this approach only requires a short observation window, in principle, it is applicable to non-stationary or even unstable systems, provided forces do not diverge during the observation time.
@@
The remainder of this paper is structured as follows. In Sec. II, we introduce and briefly discuss the model, that is, the stochastic delay differential equations that we use in our analysis and in the proof of concept. We then compute and analyze the power spectral density (defined in Sec. III.1) for a system subject to linear time-delayed feedback in Sec. III.2. After identifying the signature of time delay in linear time-delayed feedback systems, this signature serves as the basis for our general PSD-based approach to the inference of time delay. Next, after preliminary remarks (Sec. IV.1) about the treatment of non-stationary time series, we present numerical results for linear (Sec. IV.2) and specific nonlinear feedback forces (Secs. IV.2-IV.3) before demonstrating PSD-based inference for fully nonlinear systems in Sec. IV.4. Subsequently, we move on to the perturbation-based approach to infer the time delay, where we first discuss the characteristic response of a system with time delay to an external perturbation. To this end, in Sec. V.1 we first study the response of a deterministic linear system analytically using the method of steps and show that the response persists in stochastic systems and also carries over to nonlinear systems. We then discuss in Sec. V.2 how, using this response, a convolutional neural network (CNN) can be employed to classify such perturbed trajectories of time-delayed stochastic systems. Subsequently, in Sec. VI we present results for time series classification with respect to the delay time by a suitably trained neural network, demonstrating the viability and accuracy of the perturbation based approach. Again, we consider linear feedback forces in Sec. VI.1 as well as more general nonlinear feedback forces (Secs. VI.2-VI.3) in our proof of concept. This demonstrates the versatility of the neural network-based classification of time series from delayed stochastic systems. Finally, we present conclusions and an outlook to potential applications and future research directions in Sec. VII. Details about the simulation method, the inference procedures, data analysis, as well as analytical details and supplementary figures are given in Appendices.
II Model and basic equations
@@
To introduce and illustrate our approach, we consider a one-dimensional system of a simple Brownian particle subject to time-delayed feedback (dFB). The dynamics of the Brownian particle is governed by the stochastic delay differential equation (SDDE)
@@
The term $\xi(t)$ represents Gaussian white noise, with zero mean and correlation function $\langle\xi(t)\xi(t^{\prime})\rangle=2\gamma k_{B}T\delta(t-t^{\prime})$, where $k_{B}T$ (with $k_{B}$ being the Boltzmann constant and $T$ the temperature) is the thermal energy. The time-delayed feedback force, $F_{\mathrm{dFB}}[x(t),x(t-\tau)]$, depends on the instantaneous position of the particle $x(t)$ and the earlier position $x(t-\tau)$. (Here $\tau$ represents the delay time.) In principle, the SDDE can additionally include the force $F_{\mathrm{inst}}[x(t)]$ that depends only on the instantaneous position $x(t)$. In what follows, we take the particle diameter $\sigma$ as a length scale, and the Brownian time $\tau_{\mathrm{B}}\equiv\sigma^{2}/D$ as a time scale (with the bare diffusion constant $D=k_{B}T/\gamma$).
@@ II.1 Linear feedback force
For our analytical considerations, in the absence of additional instantaneous forces, we focus on a linear time-delayed feedback force $F_{\mathrm{dFB}}$ of the form
@@
The corresponding delayed feedback potential, $V_{\mathrm{quad}}$, can most intuitively be understood in the limit $k_{a}=k_{b}$, where it reduces to a harmonic potential centered at the earlier position $x(t-\tau)$. When $k_{a}\neq k_{b}$, the feedback force can be understood as competing instantaneous and delayed forces, where the dominant force is determined by the ratio of parameters $k_{a}$ and $k_{b}$. Systems subject to such linear feedback forces have been studied extensively both, in deterministic as well as stochastic systems. In the context of deterministic dynamical systems, stability analysis is one of the central points of interest (see e.g., [45, 46]) and it has, for example, been shown in an early application that time-delayed feedback can be used to stabilize otherwise unstable orbits [2, 3]. On the other hand, stability is also a central feature of stochastic systems and typically connected to stationary or steady states. For the present class of feedback forces, stability has been assessed for different parameter regimes [47] as well as in terms of extensions of the model  [48]. Here, we distinguish two cases in terms of stability. Firstly, we consider the case when $0<k_{b}\leq k_{a}$, which we refer to as attractive feedback. Attractive feedback is marginally stable for $k_{a}=k_{b}$ (attractive harmonic feedback) [11, 49], implying a constant mean position but a growing positional variance and, thus, a non-stationary time series in the language of time series analysis. Otherwise, it is stable [47, 48], i.e., both positional mean and variance are constant in the long time limit.
Secondly, the case when $k_{a}<0$ and $k_{b}<0$ we call repulsive feedback. Repulsive feedback exhibits a more complex stability behavior. In the marginally stable case, when $k_{a}=k_{b}$ (repulsive harmonic feedback), there exists a threshold value $k_{\mathrm{thresh}}(\tau)\equiv\gamma/\tau$ in the parameter space (with respect to $|k_{a}|$ and $|k_{b}|$) [50] above which the feedback forces diverge (as expected for a force derived from an inverse parabola). Below this threshold, when $|k_{a}|\leq|k_{b}|<k_{\mathrm{thresh}}(\tau)$, the system exhibits enhanced diffusion [50] or even transient super-diffusive behavior [50, 49], depending on parameter combinations and the distance from the threshold. In the present article, we study the repulsive system only for parameter combinations below $k_{\rm thresh}$.
@@ II.2 Nonlinear feedback forces
We first consider Eq. (1) with the feedback force  (4) in the absence of a heat bath (i.e., temperature $T=0$). Above a threshold in the parameter space $a\tau/(\gamma b^{2})=1$ [50], the described particle moves at a constant velocity $v_{\infty}$ in the long time limit [50] given a non-zero initial history function.
@@
Linearizations of the repulsive Gaussian feedback force (4) (leading to variants of the two marginally stable linear feedback force systems discussed above [Sec. II.1, Eq. (2) with $k_{a}=k_{b}$]) are possible for two cases (when $b$ is fixed). First, for small displacements (and $a\tau/(\gamma b^{2})<1$) we recover an repulsive harmonic feedback force [50]. Second, for the case where the particle exhibits persistent motion [50] (see Appendix B.4.2) the linearization can be understood as a co-moving attractive harmonic feedback trap.

### v2

# Inference of a time delay in stochastic systems

@@ Abstract
Time delay is ubiquitous in many experimental and real-world situations. However, from measurable data it is often unclear whether time delay plays a significant role in the dynamics, and if it does, how long the time lag really is. This would be invaluable knowledge when analyzing and modeling such systems. Hitherto, no universal method is available by which the time delay can be inferred. To address this problem, we propose and demonstrate two different methods to infer time delay in overdamped Langevin systems with delayed feedback. In the first part, we focus on the power spectral density based on the positional data and use a characteristic signature of the time delay to infer the delay time. In limiting cases, we establish a direct relation of the observations made for nonlinear time-delayed feedback forces to analytical results obtained for the linear system. In other situations despite the absence of this direct relation, the characteristic signature remains and can be exploited by a semiautomatic method to infer the delay time. Furthermore, it may not always desirable or possible to observe a system for a long time to infer dependencies and parameters. Thus, in the second part, we propose a probing method combined with a neural network to infer the delay time, which requires only short observation time series. These proposed methods for inferring time delays in stochastic systems can be valuable tools for gaining deeper insight into the role of delay across a wide range of applications – from the behavior of individual colloidal particles under feedback control to emergent collective phenomena such as flocking and swarming.
@@ I Introduction
Time delay can have a substantial impact on a wide range of experimental and real-world scenarios [1]. The presence of time delay makes the dynamics of a system dependent on the system’s history itself. In some cases, it is introduced deliberately, whereas in other situations its presence is unavoidable. A prominent example for deliberate use of delay is Pyragas time-delayed feedback for the stabilization of otherwise unstable orbits [2, 3], later applied to stabilization of satellite motion [4]. In experiments involving feedback control, researchers have explicitly considered time delay in design and modeling  [5, 6, 7]. In colloidal systems exhibiting stochastic dynamics, time-delayed feedback control is employed to modify the transport properties [8, 9]. Moreover, delay has been shown to have tremendous effect on thermodynamic observables  [10, 11, 12, 13, 14], as demonstrated in recent experiments [7, 15].
@@
Another area where delay is important and, at the same time difficult to control, are systems with finite-time information propagation. Typical examples include communication lags while steering far objects such as satellites, the reaction time humans experience when reacting to an external stimulus, or the time lag in experiments between processing image data and applying the desired steering or control impulse [16]. Moreover, time delay may have to be taken into consideration in complex systems, such as in swarms [17], in financial markets [18, 19], in robotic systems [20], and in traffic flow [21, 22]. Indeed, it has recently been shown in theoretical studies as well as simulations that the introduction of time delay and/or time-delayed feedback can induce intriguing emergent collective behavior. Examples range from polar active matter with delayed interactions [23, 24], and Gaussian time-delayed feedback applied to Brownian particles [25, 26, 27], to ecological systems [28], and oscillator networks [29, 30].
Stochastic systems with time delay belong to the class of non-Markovian systems since the dynamics depends on the systems state at an earlier time, i.e., it is history dependent. Fundamentally, the time delay may be rationalized as a consequence of hidden degrees of freedom. From a Markovian embedding perspective [31], for a system with discrete time delay, an infinite number of degrees of freedom have been projected out [32, 33]. Further, in the context of generalized Langevin equations, discrete time delay systems can be understood as systems with a $\delta$-shaped memory kernel. Notice that this memory kernel, however, is not related to the environmental noise correlation [32].
@@
However, in many experimental and real-world situations, it is not a priori clear whether time delay plays a significant role [34, 16] or if it is present at all. Therefore, it is of fundamental importance to investigate the presence of time delay based on observable quantities to accurately design and understand physical systems.
In this study, we focus on systems with a specific delay time rather than multiple delays or distributed delay [35]. Of course, assuming a single discrete time delay in a real system is a rather strong assumption. Broadly speaking, systems with several time delays, distributed delay or a more general memory kernel have been studied. However, there are numerous systems where the discrete time delay assumption is well justified. In laser systems [36, 37] and quantum-optical systems [38, 39], time delay originates from light traveling through the optical setup at finite speed. Further examples include complex networks with finite signal propagation speeds (see Refs. [1, 40] and references therein on the topics of neural and laser networks). In other cases, one can expect a ‘typical’ time delay as part of a peaked distribution, e.g., when considering finite processing and reaction times. Therefore, gaining information about the presence of time delay as well as the ‘typical’ or average time delay is of high value.
@@
In recent years, ideas of parameter inference for dynamical systems and equation discovery (for ordinary as well as partial differential equations) from (noisy) data have received considerable interest. Methods range from Bayesian inference [41] and regression-based machine learning [42] to the application of neural network-based techniques [43]. For some types of system, methods for identifying governing equations [44, 45] and inferring parameters for deterministic and linear stochastic delay differential equations (DDE and SDDE) [46, 47], specifically time delay in the underlying dynamics from time series data [48, 49, 50] exist. However, those methods are typically designed for application to deterministic systems or noisy measurement data [49]. For fully stochastic (non)linear systems, on the other hand, to the best of our knowledge, no universal approach for the inference of time delay exists. Motivated by this gap, we here propose and demonstrate, for the first time, two approaches to infer a time delay for Brownian systems.
@@
The first approach is based on the power spectral density (PSD) of stochastic time series, which is often used to analyze experimental data [51, 52, 53] including systems with anomalous diffusion [54]. Here, we identify a characteristic signature of time delay, namely oscillations with a delay-specific period, and demonstrate in a proof of concept how this signature can be exploited to infer the delay time. To do this, we start from Brownian systems subject to linear time-delayed feedback as well as specific delayed nonlinear forces including respective limiting cases and then consider highly nonlinear time-delayed feedback. We thereby demonstrate that the characteristic signature carries over from systems subject to linear feedback forces to nonlinear systems and, thus, that the PSD-based inference approach is agnostic, i.e., it does not rely on knowledge of the functional form of the underlying stochastic equations of motion.
@@
The PSD approach, however, has a drawback. A large number (typically $N\gtrsim 100$) of relatively long [several Brownian times $\tau_{B}$ (defined in Sec. II)] trajectories is required to infer the time delay with sufficient precision. In real-world applications, having to record a large number of long trajectories may (depending on the scenario) not be desirable or not even possible.
@@
Given the potential difficulties, we propose a powerful alternative approach employing a neural network. To establish the concept, we utilize the fact that systems that involve time delay, rather than instantaneously responding to an external stimulus, react with some delay. Therefore, by applying a suitable perturbation and observing the dynamical response, we can obtain information about the underlying delay time of the system. In order to classify the system’s response, based on trajectory time series data, we make use of a convolutional neural network [55, 56]. Such a network can be trained to output the most likely time delay for one or even several different functional forms of the underlying delayed and instantaneous forces. Compared to the PSD approach, the main advantage of this approach is that the observation window (i.e., the trajectory length) only has to be of the order of the time delay itself, and relatively few trajectories (typically $N\sim 25$) are sufficient for precise inference. In combination with state-of-the-art computer systems, this brings the inference of time delay in real time within the realm of possibilities.
@@
The remainder of this paper is structured as follows. In Sec. II, we introduce and briefly discuss the model, that is, the stochastic delay differential equations that we use in our analysis and in the proof of concept. We then compute and analyze the power spectral density (defined in Sec. III.1) for a system subject to linear time-delayed feedback in Sec. III.2. After identifying the signature of time delay in linear time-delayed feedback systems, this signature serves as the basis for our general PSD-based approach to the inference of time delay. Next, after preliminary remarks (Sec. IV.1) about the treatment of non-stationary time series, we present numerical results for linear (Sec. IV.2) and specific nonlinear feedback forces (Sec. IV.3) before demonstrating PSD-based inference for fully nonlinear systems in Sec. IV.4. Subsequently, we move on to the perturbation-based approach to infer the time delay, where we first discuss the characteristic response of a system with time delay to an external perturbation. To this end, in Sec. V.1 we study the response of a deterministic linear system analytically using the method of steps and show that the response persists in stochastic systems and also carries over to nonlinear systems. We then discuss in Sec. V.2 how, using this response, a convolutional neural network (CNN) can be employed to classify such perturbed trajectories of time-delayed stochastic systems. Subsequently, in Sec. VI we present results for time series classification with respect to the delay time by a suitably trained neural network, demonstrating the viability and accuracy of the perturbation based approach. Again, we consider linear feedback forces in Sec. VI.1 as well as more general nonlinear feedback forces (Secs. VI.2-VI.3) in our proof of concept. This demonstrates the versatility of the neural network-based classification of time series from delayed stochastic systems. Finally, we present conclusions and an outlook to potential applications and future research directions in Sec. VII. Details about the simulation method, the inference procedures, data analysis, as well as analytical details, supplementary figures and a brief discussion of colored noise systems with corresponding numerical data are given in Appendices.
II Model
@@
To set the stage, we consider a one-dimensional system of a Brownian particle subject to time-delayed feedback (dFB). The dynamics of the Brownian particle is governed by the stochastic delay differential equation (SDDE)
@@
The term $\xi(t)$ represents Gaussian white noise, with zero mean and correlation function $\langle\xi(t)\xi(t^{\prime})\rangle=2\gamma k_{B}T\delta(t-t^{\prime})$, where $k_{B}T$ (with $k_{B}$ being the Boltzmann constant and $T$ the temperature) is the thermal energy. (We here focus on Gaussian white noise for simplicity. However, Appendix D discusses the applicability and numerical results based on the methods introduced below for systems with exponentially correlated noise.) The time-delayed feedback force, $F_{\mathrm{dFB}}[x(t),x(t-\tau)]$, depends on the instantaneous position of the particle $x(t)$ and the earlier position $x(t-\tau)$, where $\tau$ represents the delay time. Instead of an initial condition, as it is required for non-delayed systems, the SDDE is complemented with an initial history function $\Phi(t)$ containing the particle’s position for $t\in[-\tau,0]$. In principle, the SDDE can additionally include the force $F_{\mathrm{inst}}[x(t)]$ that depends only on the instantaneous position $x(t)$. In what follows, we take the particle diameter $\sigma$ as a length scale, and the Brownian time $\tau_{\mathrm{B}}\equiv\sigma^{2}/D$ as a time scale (with the bare diffusion constant $D=k_{B}T/\gamma$).
@@ II.1 Linear feedback force
For our analytical considerations we focus on a linear time-delayed feedback force $F_{\mathrm{dFB}}$ of the form
@@
The corresponding delayed feedback potential, $V_{\mathrm{quad}}$, can most intuitively be understood in the limit $k_{a}=k_{b}$, where it reduces to a harmonic potential centered at the earlier position $x(t-\tau)$. When $k_{a}\neq k_{b}$, the feedback force can be understood as competing instantaneous and delayed forces, where the dominant force is determined by the ratio of parameters $k_{a}$ and $k_{b}$. Systems subject to such linear feedback forces have been studied extensively both, in deterministic [2, 57, 4, 58] as well as stochastic systems [57, 59, 12, 13, 10]. In the context of deterministic dynamical systems, stability analysis is one of the central points of interest (see e.g., [60, 58]) and it has, for example, been shown in an early application that time-delayed feedback can be used to stabilize otherwise unstable orbits [2, 3]. On the other hand, stability is also a central feature of stochastic systems and typically connected to stationary or steady states. For the present class of feedback forces, stability has been assessed for different parameter regimes [57] as well as in terms of extensions of the model  [35]. Here, we distinguish two cases in terms of stability. Firstly, we consider the case when $0<k_{b}\leq k_{a}$, which we refer to as attractive feedback. Specifically, attractive feedback is marginally stable for $k_{a}=k_{b}$ (attractive harmonic feedback) [11, 61], implying a constant mean position but a growing positional variance and, thus, a non-stationary time series in the language of time series analysis. Otherwise, it is stable [57, 35], i.e., both positional mean and variance are constant in the long time limit.
Secondly, the case when $k_{a}<0$ and $k_{b}<0$ we call repulsive feedback. Repulsive feedback exhibits a more complex stability behavior. In the marginally stable case, when $k_{a}=k_{b}$ (repulsive harmonic feedback), there exists a threshold value $k_{\mathrm{thresh}}(\tau)\equiv\gamma/\tau$ in the parameter space (with respect to $|k_{a}|$ and $|k_{b}|$) [62] above which the feedback forces diverge (as expected for a force derived from an inverse parabola). Below this threshold, when $|k_{a}|=|k_{b}|<k_{\mathrm{thresh}}(\tau)$, the system exhibits enhanced diffusion [62] or even transient super-diffusive behavior [62, 61], depending on parameter combinations and the distance from the threshold. In the present article, we study the repulsive system only for parameter combinations below $k_{\rm thresh}$. Furthermore, when $|k_{a}|<|k_{b}|<k_{\mathrm{thresh}}(\tau)$ and $\tau<\arccos(-k_{a}/k_{b})/\sqrt{k_{b}^{2}-k_{a}^{2}}$ the system assumes a stationary state [57]. These conditions are satisfied in our considerations for repulsive feedback.
@@ II.2 Nonlinear feedback forces
To understand the impact of $F_{\mathrm{Gauss}}$, we first consider Eq. (1) with the feedback force (4) in the absence of a heat bath (i.e., temperature $T=0$). Above a threshold in the parameter space, i.e., $a\tau/(\gamma b^{2})>1$  [62], the described particle moves at a constant velocity $v_{\infty}$ in the long time limit [62] given a non-zero initial history function.
@@
Linearizations of the repulsive Gaussian feedback force (4) (leading to variants of the two marginally stable linear feedback force systems discussed above [Sec. II.1, Eq. (2) with $k_{a}=k_{b}$]) are possible for two cases (when $b$ is fixed). First, for small displacements (and $a\tau/(\gamma b^{2})<1$) we recover a repulsive harmonic feedback force [62] (see Appendix A.1.1). Second, for the case where the particle exhibits persistent motion [62] (see Appendix A.1.2) the linearization can be understood as a co-moving attractive harmonic feedback trap.